# WhiteFox: White-Box Compiler Fuzzing Empowered by Large Language Models

### v2

@@ Abstract.
Compiler correctness is crucial, as miscompilation can falsify program behaviors, leading to serious consequences over the software supply chain. In the literature, fuzzing has been extensively studied to uncover compiler defects. However, compiler fuzzing remains challenging: Existing arts focus on black- and grey-box fuzzing, which generates test programs without sufficient understanding of internal compiler behaviors. As such, they often fail to construct test programs to exercise intricate optimizations. Meanwhile, traditional white-box techniques, such as symbolic execution, are computationally inapplicable to the giant codebase of compiler systems. Recent advances demonstrate that Large Language Models (LLMs) excel in code generation/understanding tasks and even have achieved state-of-the-art performance in black-box fuzzing. Nonetheless, guiding LLMs with compiler source-code information remains a missing piece of research in compiler testing.
@@ 2. Introduction
In the literature, researchers have proposed various fuzzing techniques to incorporate the knowledge about the system under test (SUT) during test generation (Even-Mendoza et al., 2023; Chen et al., 2021b). They are generally classified into three categories according to the extent of SUT knowledge visible to the fuzzer: black-box (Liu et al., 2023b; Yang et al., 2011; Livinskii et al., 2020), grey-box (libFuzzer, 2023; Even-Mendoza et al., 2023; Chen et al., 2021b), and white-box fuzzing (King, 1976; Sen, 2007). Black-box fuzzing has zero knowledge about the internal workings of the SUT and simply considers system input/interface information. Consequently, the inputs are generated without complying with the intended structures or behaviors of the SUT. In contrast, white-box fuzzing techniques, by inspecting the source code of the SUT, aim to synthesize test cases to exhaustively explore all possible code paths. Grey-box fuzzing lies between black- and white-box fuzzing. By leveraging limited program information of the SUT (e.g., code coverage), grey-box fuzzers attempt to efficiently produce tests that are more likely to exercise new program behaviors. While in theory, we can apply all such approaches for compiler fuzzing, each approach grapples with its challenges and limitations owing to the immense complexity and scale of modern compiler systems. For instance, the widely-used LLVM (Lattner and Adve, 2004) compiler is implemented with 14M lines of code, and the popular DL compiler, TensorFlow (TensorFlow, 2023), has 3.5M lines.
@@
Challenges. Black-box fuzzing, without knowledge of the internal workings, struggles due to the intricate conditions required to trigger optimizations. Simply generating random inputs, without any guidance, often proves impractical for reaching the deep corners of optimization logic. For example, a recent study (Chen et al., 2021b) shows that the black-box fuzzer for C compilers, Csmith (Yang et al., 2011), which produces test-cases through grammar-based generation, is significantly less effective than coverage-guided fuzzing. Grey-box fuzzing, though better informed by source code instrumentation to achieve higher code coverage than its black-box counterpart, frequently falls short of fully understanding the nuanced criteria required to trigger particular optimizations. This shortcoming stems from the fact that compiler optimizations typically hinge on meeting precise and strict conditions. Vanilla coverage-driven strategies might not navigate these specific states effectively. Moreover, grey-box compiler fuzzing (Chen et al., 2021b; Even-Mendoza et al., 2023) even fail to generate semantically correct inputs, leading to the discovery of mostly front-end crash bugs. On the other hand, traditional white-box fuzzing, which relies on strict analysis of the SUT source code, becomes daunting with modern compilers. The sheer complexity of modern optimization techniques, combined with the vast landscape of programming paradigms and hardware targets, makes modeling all behaviors an uphill task. For instance, symbolic execution (King, 1976) executes a program by using symbolic variables in place of concrete values, enabling the systematic exploration of every potential execution path. However, when applied to compiler systems, it becomes infeasible to designate every variable as symbolic. Even if such a feat were achievable, the million-line scale of compiler codebase inevitably leads to path explosion, rendering the approach highly challenging.
@@
Insight. Can we scale white-box fuzzing to fully test optimizations for any compilers? We address this question based on the insight that Large Language Models (LLMs) (OpenAI, 2023b, a; Li et al., 2023a; Chen et al., 2021a; Yue Wang and Hoi, 2021; Feng et al., 2020) are pre-trained on a vast array of code spanning various programming languages. This broad foundation enables them to excel in comprehending and generating code across diverse languages (Bubeck et al., 2023).Therefore, for the permute_linear_fusion optimization, different from typical white-box fuzzing, we can leverage LLMs to summarize the requirements for the models that could trigger it based on the source code information, as highlighted in the yellow text box of Figure 1. Subsequently, we can leverage the generated requirement description to further prompt LLMs to create the corresponding inputs, which are PyTorch models in this case. In our experiments, the generated tests indeed triggered the permute_linear_fusion optimization and even detected a previously unknown bug in it! Notably, this bug was confirmed by developers and labeled as high-priority.
@@
  • Approach. While our approach is general and applicable to various compiler systems, we implement WhiteFox as a practical fuzzer for the three most popular DL compilers, PyTorch Inductor, TensorFlow-XLA, and TensorFlow Lite. We utilize GPT4 as an analysis agent to summarize the requirements based on the source code, and StarCoder as a generation agent to create diverse test inputs. Our artifact is available at https://github.com/ise-uiuc/WhiteFox.
@@ 3.2. LLM-based Testing
Beyond unit test generation, applying LLMs to fuzz software systems is another emerging trend (Xia et al., 2023; Deng et al., 2023b, a; Sun et al., 2023). For example, TitanFuzz (Deng et al., 2023a) demonstrates for the first time that modern LLMs can be directly leveraged to perform both grammar- and mutation-based fuzzing of real-world systems. Very recently, Fuzz4All (Xia et al., 2023) has also demonstrated that LLMs can serve as the universal fuzzers for various types of software systems. However, these approaches often neglect white-box information, specifically the source code implementation of the software under test. Consequently, they tend to concentrate on the compiler’s front-end rather than the intricate logic within the middle and back-end components. In contrast, WhiteFox is the first white-box compiler fuzzing approach, which leverages low-level optimization implementations for guided fuzzing, and has been demonstrated to outperform the recent TitanFuzz (§ 7.1).
@@ 4.3. Feedback Loop
Figure 5. Prompting for test generation with feedback.
@@ 6.2. Experimental Setup
Fuzzing budget. Our default setting is to generate a total of 1000 tests for each optimization in 100 iterations. In each iteration, WhiteFox by default generates a batch of 10 tests based on optimization triggering feedback from previous iterations. If the optimization was triggered in previous iterations, WhiteFox picks three triggering inputs used as few-shot examples in the feedback-guided prompt (Section  4.3, Figure  5) for the following iterations. Otherwise, WhiteFox uses the default few-shot prompt (Section  4.2, Figure  4) to generate tests.
@@ 6.3. Metrics
To further show the effectiveness of every component, we also use code coverage (Deng et al., 2023a; Liu et al., 2023b; Even-Mendoza et al., 2023; Klees et al., 2018) as a metric. Specifically, we report line coverage in the source languages where the optimizations are implemented: Python for PyTorch Inductor, and C++ for both TensorFlow Lite and TensorFlow-XLA. Following previous work  (Wei et al., 2022; Gu et al., 2022; Yang et al., 2023), we measure line coverage using Coverage.py (cov, 2022) for Python and GCOV (gco, 2022) for C++.
@@ 7.1. Comparison with Prior Work
Additionally, Table 3 compares WhiteFox with the baselines over a 24-hour testing period, a common setting for fuzzing approaches (Klees et al., 2018). WhiteFox performs the best on all three subjects, substantially outperforming others on PyTorch Inductor and TensorFlow Lite. In terms of code coverage, WhiteFox covers more lines than the baselines in PyTorch Inductor and TensorFlow-XLA by up to 19.9%. For TensorFlow Lite, WhiteFox performs slightly worse than TitanFuzz (5.9%). This may be attributed to the limited number of optimizations (13) in TensorFlow Lite, which inherently restricts WhiteFox’s code coverage exploration, as WhiteFox does not have much white-box information (i.e., optimization implementation) to guide the generation. Meanwhile, please kindly note that code coverage is just a proxy indicator and does not always correlate strongly with bug finding abilities for complicated systems (Inozemtseva and Holmes, 2014; Su et al., 2021). Despite slightly lower code coverage on TensorFlow Lite, WhiteFox still performs better on the ultimate goal, bug finding (detailed in Section 7.3). Overall, these results demonstrate the effectiveness of WhiteFox in generating test cases to cover not only optimizations but also various compiler behaviors.
@@ 7.3. Bug Finding
Table 6 presents the bug-finding results of WhiteFox over PyTorch, TensorFlow-XLA, and TensorFlow Lite. To date, WhiteFox has detected 101 bugs for them. Of these, 92 are confirmed as previously unknown, and 70 are already fixed. For PyTorch, we conducted bug detection in both the evaluation and the latest versions upon the request of the PyTorch official team (will be discussed in § 8.1). Thus, among the 79 identified bugs in the PyTorch Inductor, 14 are found in its most recent version. Remarkably, 10 (12.7%) of the PyTorch bugs have been labeled with high priority.
@@ 8.3. Limitations and Future Work
Another possible future direction is to use traditional fuzzing approaches as external tools for efficient test generation. Given the higher computational cost of invoking smaller LLMs compared to traditional techniques, this strategy could improve performance. For instance, we could utilize WhiteFox to summarize optimization triggering generation or mutation rules, which could then guide input generation for a traditional fuzzing framework such as NNSmith (Liu et al., 2023b).
@@ Acknowledgements.
We thank Steven Chunqiu Xia for providing the help and resources to run some experiments. This work was partially supported by NSF grant CCF-2131943 and Kwai Inc. This project is supported, in part, by funding from Two Sigma Investments, LP. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Two Sigma Investments, LP.

### v3

@@ Abstract.
Compiler correctness is crucial, as miscompilation can falsify program behaviors, leading to serious consequences over the software supply chain. In the literature, fuzzing has been extensively studied to uncover compiler defects. However, compiler fuzzing remains challenging: Existing solutions focus on black- and grey-box fuzzing, which generates test programs without sufficient understanding of internal compiler behaviors. As such, they often fail to construct test programs to exercise intricate optimizations. Meanwhile, traditional white-box techniques, such as symbolic execution, are computationally inapplicable to the giant codebase of compiler systems. Recent advances demonstrate that Large Language Models (LLMs) excel in code generation/understanding tasks and even have achieved state-of-the-art performance in black-box fuzzing. Nonetheless, guiding LLMs with compiler source-code information remains a missing piece of research in compiler testing.
@@ 1. Introduction
In the literature, researchers have proposed various fuzzing techniques to incorporate the knowledge about the system under test (SUT) during test generation (Even-Mendoza et al., 2023; Chen et al., 2021b). They are generally classified into three categories according to the extent of SUT knowledge visible to the fuzzer: black-box (Liu et al., 2023b; Yang et al., 2011; Livinskii et al., 2020), grey-box (libFuzzer, 2023; Even-Mendoza et al., 2023; Chen et al., 2021b), and white-box fuzzing (King, 1976; Sen, 2007). Black-box fuzzing has zero knowledge about the internal workings of the SUT and simply considers system input/interface information. Consequently, the inputs are generated without complying with the intended structures or behaviors of the SUT. In contrast, white-box fuzzing techniques, by inspecting the source code of the SUT, aim to synthesize test cases to exhaustively explore all possible code paths. Grey-box fuzzing lies between black- and white-box fuzzing. By leveraging limited program information of the SUT (e.g., code coverage), grey-box fuzzers attempt to efficiently produce tests that are more likely to exercise new program behaviors. While in theory, we can apply all such approaches for compiler fuzzing, each approach grapples with its challenges and limitations owing to the immense complexity/scale of modern compiler systems. For instance, the widely-used LLVM (Lattner and Adve, 2004) compiler is implemented with 14M lines of code, and the popular DL compiler, TensorFlow (TensorFlow, 2023), has 3.5M lines.
@@
Challenges. Black-box fuzzing, without knowledge of the internal workings, struggles due to the intricate conditions required to trigger optimizations. Simply generating random inputs, without any guidance, often proves impractical for reaching the deep corners of optimization logic. For example, a recent study (Chen et al., 2021b) shows that the black-box fuzzer for C compilers, Csmith (Yang et al., 2011), which produces test-cases through grammar-based generation, can be significantly less effective than coverage-guided fuzzing. Grey-box fuzzing, though better informed by source code instrumentation to achieve higher code coverage than its black-box counterpart, frequently falls short of fully understanding the nuanced criteria required to trigger particular optimizations. This shortcoming stems from the fact that compiler optimizations typically hinge on meeting precise and strict conditions. Vanilla coverage-driven strategies might not navigate these specific states effectively. Moreover, grey-box compiler fuzzing (Chen et al., 2021b; Even-Mendoza et al., 2023) even fail to generate semantically correct inputs, leading to the discovery of mostly front-end crash bugs. On the other hand, traditional white-box fuzzing, which relies on strict analysis of the SUT source code, becomes daunting with modern compilers. The sheer complexity of modern optimization techniques, combined with the vast landscape of programming paradigms and hardware targets, makes modeling all behaviors an uphill task. For instance, symbolic execution (King, 1976) executes a program by using symbolic variables in place of concrete values, enabling the systematic exploration of every potential execution path. However, when applied to compiler systems, it becomes infeasible to designate every variable as symbolic. Even if such a feat were achievable, the million-line scale of compiler codebase inevitably leads to path explosion, rendering the approach highly challenging.
@@
Insight. Can we scale white-box fuzzing to fully test optimizations for any compilers? We address this question based on the insight that Large Language Models (LLMs) (OpenAI, 2023b, a; Li et al., 2023a; Chen et al., 2021a; Yue Wang and Hoi, 2021; Feng et al., 2020; Wei et al., 2024) are pre-trained on a vast array of code spanning various programming languages. This broad foundation enables them to excel in comprehending and generating code across diverse languages (Bubeck et al., 2023).Therefore, for the permute_linear_fusion optimization, different from typical white-box fuzzing, we can leverage LLMs to summarize the requirements for the models that could trigger it based on the source code information, as highlighted in the yellow text box of Figure 1. Subsequently, we can leverage the generated requirement description to further prompt LLMs to create the corresponding inputs, which are PyTorch models in this case. In our experiments, the generated tests indeed triggered the permute_linear_fusion optimization and even detected a previously unknown bug in it! Notably, this bug was confirmed by developers and labeled as high-priority.
@@
  • Approach. While our approach is general and applicable to various compiler systems, we implement WhiteFox as a practical fuzzer for the three most popular DL compilers, PyTorch Inductor, TensorFlow-XLA, and TensorFlow Lite. We utilize GPT4 (OpenAI, 2023b) as an analysis agent to summarize the requirements based on the source code, and StarCoder (Li et al., 2023a) as a generation agent to create diverse test inputs. Our artifact is available at https://github.com/ise-uiuc/WhiteFox.
@@ 2.2. LLM-based Testing
Beyond unit test generation, applying LLMs to fuzz software systems is another emerging trend (Xia et al., 2023; Deng et al., 2023b, a; Sun et al., 2023; Huang et al., 2024; Yang et al., 2023b; Meng et al., 2024; Ou et al., 2024). For example, TitanFuzz (Deng et al., 2023a) demonstrates for the first time that modern LLMs can be directly leveraged to perform both grammar- and mutation-based fuzzing of real-world systems. Recently, Fuzz4All (Xia et al., 2023) has also demonstrated that LLMs can serve as the universal fuzzers for various types of software systems. However, these approaches often neglect white-box information, specifically the source code implementation of the software under test. Consequently, they tend to concentrate on the compiler’s front-end rather than the intricate logic within the middle and back-end components. In contrast, WhiteFox is the first white-box compiler fuzzing approach, which leverages low-level optimization implementations for guided fuzzing, and has been demonstrated to outperform the recent TitanFuzz (§ 6.1).
@@ 19             $\beta_{new}\leftarrow\beta_{avg}$
Figure 5. Prompting for test generation with feedback.
@@ 5.2. Experimental Setup
Fuzzing budget. Our default setting is to generate a total of 1000 tests for each optimization in 100 iterations. In each iteration, WhiteFox by default generates a batch of 10 tests based on optimization triggering feedback from previous iterations. If the optimization was triggered in previous iterations, WhiteFox picks three triggering inputs used as few-shot examples in the feedback-guided prompt (Section  3.3, Figure  5) for the following iterations. Otherwise, WhiteFox uses the default few-shot prompt (Section  3.2, Figure  4) to generate tests.
@@ 5.3. Metrics
To further show the effectiveness of every component, we also use code coverage (Deng et al., 2023a; Liu et al., 2023b; Even-Mendoza et al., 2023; Klees et al., 2018) as a metric. Specifically, we report line coverage in the source languages where the optimizations are implemented: Python for PyTorch Inductor, and C++ for both TensorFlow Lite and TensorFlow-XLA. Following previous work  (Wei et al., 2022; Gu et al., 2022; Yang et al., 2023a), we measure line coverage using Coverage.py (cov, 2022) for Python and GCOV (gco, 2022) for C++.
@@ 6.1. Comparison with Prior Work
Additionally, Table 3 compares WhiteFox with the baselines over a 24-hour testing period, a common setting for fuzzing studies (Klees et al., 2018). WhiteFox performs the best on the number of triggered optimizations, substantially outperforming others on PyTorch Inductor and TensorFlow Lite. In terms of code coverage, WhiteFox covers more lines than the baselines in PyTorch Inductor and TensorFlow-XLA by up to 19.9%. For TensorFlow Lite, WhiteFox performs slightly worse than TitanFuzz (5.9%). This may be attributed to the limited number of optimizations (13) in TensorFlow Lite, which inherently restricts WhiteFox’s code coverage exploration, as WhiteFox does not have much white-box information (i.e., optimization implementation) to guide the generation. Meanwhile, please kindly note that code coverage is just a proxy indicator and does not always correlate strongly with bug finding abilities for complicated systems (Inozemtseva and Holmes, 2014; Su et al., 2021). Despite slightly lower code coverage on TensorFlow Lite, WhiteFox still performs better on the ultimate goal, bug finding (detailed in Section 6.3). Overall, these results demonstrate the effectiveness of WhiteFox in generating test cases to cover not only optimizations but also various compiler behaviors.
@@ 6.3. Bug Finding
Table 6 presents the bug-finding results of WhiteFox over PyTorch, TensorFlow-XLA, and TensorFlow Lite. To date, WhiteFox has detected 101 bugs for them. Of these, 92 are confirmed as previously unknown, and 70 are already fixed. For PyTorch, we also conducted bug detection in the latest version upon the request of the PyTorch team (will be discussed in § 7.1). Thus, among the 79 identified bugs in the PyTorch Inductor, 14 are found in its most recent version. Remarkably, 10 (12.7%) of the PyTorch bugs have been labeled with high priority.
@@ 7.3. Limitations and Future Work
Another possible future direction is to use traditional fuzzing approaches as external tools for efficient test generation. Given the higher computational cost of invoking smaller LLMs compared to traditional techniques, this strategy could improve performance. For instance, we could utilize WhiteFox to summarize optimization-triggering generation or mutation rules, which could then guide input generation for a traditional fuzzing framework such as NNSmith (Liu et al., 2023b).
@@ Acknowledgements.
We thank Chunqiu Steven Xia for providing the help and resources to run some experiments. This work was partially supported by NSF grant CCF-2131943 and Kwai Inc. This project is supported, in part, by funding from Two Sigma Investments, LP. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Two Sigma Investments, LP.